A random variable x is binomially distributed with mean 3 and variance 2. find the probability distribution of X, and also draw the graph of Distribution Function.
Solution:
as given mean=np=3-----(1) and var(x) =npq= 2--------(2)
substituting value fro 1 into 2.
2=(3)q => q=2/3

p=1-2/3 = 1/3
eq(1)=>
n(1/3)=3 => n=9
then after that what whould i do to calculate the probability distributions for x??? please guide me sir.. I shall be very grateful. Thanks!

Answer by rothauserc(4718)

The binomial distribution is a discrete distribution that approximates a normal distribution
:
The graph of a binomial distribution is a histogram
:
you have calculated p, q, n
:
use binomial probability calculator for each k from 0 to 9 and plot your histogram, the x axis will be 0 through 9 and y axis will be the associated probabilities
:
the binomial probability formula is
:
probability(k successes in n trials) = nCk * p^k * (1-p)^(n-k) where n=9, p=1/3 and nCk = n!/(k! * (n-k)!)
: